OpenStudy (anonymous):

Widget Wonders produces widgets. They have found that the cost, c(x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $16 to produce 2 widgets, $18 to produce 4 widgets, and $48 to produce 10 widgets. Find the total cost of producing 8 widgets. ( you use quadratic function formula)

OpenStudy (anonymous):

Any ideas???

OpenStudy (anonymous):

@Jamierox4ev3r @bibby @Directrix @Goten77 @mathstudent55 @Firejay5

OpenStudy (anonymous):

not at all @pielover123

OpenStudy (anonymous):

@GabbyGabsterz @Holly00d1248 @Jamielynne @Michele_Laino @leletoomuch

OpenStudy (anonymous):

c(x) = ax^2 + bx + d for some a, b, and d Substitute (2, 30), (4, 118), and (10, 766) for (x, y): 30 = a(2^2) + b(2) + d 118 = a(4^2) + b(4) + d 766 = a(10^2) + b(10) + d Now you can solve the system of equations for a, b, and d, which will give you c(x)

OpenStudy (anonymous):

Are you lost or do you follow?

OpenStudy (anonymous):

and how did you get that @pielover123

OpenStudy (anonymous):

Do you know what X=?

OpenStudy (anonymous):

8 @pielover123

OpenStudy (anonymous):

Total Cost for 8 Widgets = $320

OpenStudy (michele_laino):

why (2, 30), (4, 118), and (10, 766)? @pielover123

OpenStudy (anonymous):

My dad is helping me Honest i just wrote down what he told me to put

OpenStudy (anonymous):

cant be 320 becasue it cost 48$ to produce 10 @pielover123

OpenStudy (michele_laino):

I think this @pielover123 : 16 = a(2^2) + b(2) + d 18 = a(4^2) + b(4) + d 48 = a(10^2) + b(10) + d

OpenStudy (anonymous):

your right but once you simply that and put it in the calculator it says no solution which is the part thats confusing me @Michele_Laino

OpenStudy (michele_laino):

please wait I try to compute your solution @bdfhbfg

OpenStudy (anonymous):

alright thanks i just need to find out what a,b,and c would be @Michele_Laino

OpenStudy (michele_laino):

I got these values: a= 1/2 b= - 2 d= 18 so your function, is: \[c(x) = \frac{{{x^2}}}{2} - 2x + 18\]

OpenStudy (michele_laino):

then your answer is: \[c\left( 8 \right) = \frac{{{8^2}}}{2} - 2 \cdot 8 + 18 = ...?\]

OpenStudy (anonymous):

hold on for a second @Michele_Laino

OpenStudy (anonymous):

THANK YOUUUUUUUU YOU LITERALLY SAVED MY GRADE!. @Michele_Laino

OpenStudy (michele_laino):

thank you! :)

OpenStudy (anonymous):

appreciate you man! ! great guy @Michele_Laino

OpenStudy (michele_laino):

thanks! again!